OpenStudy (anonymous):

Hi all! How do I take the derivative of the inverse tangent of (x^2) ?

OpenStudy (anonymous):

Do you know the derivative of the inverse tangent of x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1/(1+x^2)

OpenStudy (anonymous):

so is it 1/(1+x^4) ?

OpenStudy (anonymous):

yes, almost but remeber that when taking the derivative of example sin(x^2) it isn't cos(x^2) but you also have to multiply with the derivative of x^2 so it's 2x*cos(x^2). Can you apply that to the arctan case?

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

are you talking about the chain rule?

OpenStudy (anonymous):

Precisely, when we take the derivative of f(x^2) we say x^2=u => f(x^2) = f(u) and then the derivative is D(u)*D(f(u))

OpenStudy (anonymous):

For the arctan case you have to multiply what you got with the derivative of x^2 = 2x

OpenStudy (anonymous):

ah, ok

OpenStudy (anonymous):

hmm so it is 1/(1+2x^2 *2x)

OpenStudy (anonymous):

ok, so generally we have as above according to the chain rule: D(u)*D(f(u)) and in this case f(u) = arctan(u). The derivative of arctan(u) is as you stated above = 1/(u^2+1) and the derivative of u = 2x. Which means we have D(u)*D(f(u)) = 2x*(1/((x^2)^2+1)) = 2x/(1/((x^4)+1))

OpenStudy (anonymous):

sorry I ment 2x/((x^4)+1))

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

I thought I would just take the chain rule of the bottom

OpenStudy (anonymous):

where (x^2)^2

OpenStudy (anonymous):

I see, it makes sense

OpenStudy (anonymous):

obviously I'm dealing with the entire function

OpenStudy (anonymous):

Yes, precisely

OpenStudy (anonymous):

cool, thanks!

OpenStudy (anonymous):

No problem :)